On Monday, Serena averaged 40 mph while making sales calls. On Tuesday, she drove for 2 hours more than she did on Monday and averaged 50 mph. If her total distance for the two days was 280 miles, then how many hours did she drive on Monday?: On Monday, Serena averaged 40 mph while making sales calls. On Tuesday, she drove for 2 hours more than she did on Monday and averaged 50 mph. If her total distance for the two days was 280 miles, then how many hours did she drive on Monday?
Answer by mangopeeler07(441)

x=amount of hours on Monday
coefficient=mph

40(x)+50(x+2)=280

Simplify
40+50x+100=280

Combine like terms
50x+140=280

Subtract 140 from both sides
50x=140

Divide by 50
x=2.8

She drove 2.8 hours on Monday.
Question 149146: On Monday, Serena averaged 40 mph while making sales calls. On Tuesday, she drove for 2 hours more than she did on Monday and averaged 50 mph. If her total distance for the two days was 280 miles, then how many hours did she drive on Monday?: On Monday, Serena averaged 40 mph while making sales calls. On Tuesday, she drove for 2 hours more than she did on Monday and averaged 50 mph. If her total distance for the two days was 280 miles, then how many hours did she drive on Monday?
Answer by josmiceli(1901)

For Monday:
d = r*t
d = 40t
For Tuesday
280 - d = r*t
280 - d = 50*(t + 2)
This is 2 equations with 2 unknowns
280 - 40t = 50*(t + 2)
280 - 40t = 50t + 100
90t = 180
t = 2 hrs
She drove 2 hrs on Monday
check:
d = 40t
d = 40*2
d = 80 mi
280 - d = 50*(t + 2)
280 - d = 50*4
280 - d = 200
d = 80 mi
OK